### v2

# Homogeneous potentials, Lagrange’s identity and Poisson geometry

Abstract
The Lagrange identity expresses the second derivative of the moment of inertia of a system of material points through kinetic energy and homogeneous potential energy, from which follows the Jacobi well-known result on the instability of a system of gravitating bodies. In this work, it is proven that if a Hamiltonian system satisfies the Lagrange identity, then it possesses additional tensor invariants that are not expressed through the basic invariants existing for all Hamiltonian systems. A new class of Hamiltonian systems with inhomogeneous potentials is considered, which also possess similar additional tensor invariants.
1 Introduction
Let us consider the Hamiltonian system of differential equations
$\frac{dq_{i}}{dt}=\frac{\partial H}{\partial p_{i}}\,,\qquad\frac{dp_{i}}{dt}=-\frac{\partial H}{\partial q_{i}}\,,\qquad i=1,\ldots,n,$ (1.1)
where $q=\{q_{1},\ldots,q_{n}\}$ are coordinates in configuration space $\mathbb{R}^{n}$, $p=\{p_{1},\ldots,p_{n}\}$ are the corresponding moments in phase space $T^{*}\mathbb{R}^{n}$, and $H=H(q,p)$ is the Hamiltonfunction.
If the Hamilton function
$H=T+V=\frac{1}{2m_{1}}p_{1}^{2}+\cdots+\frac{1}{2m_{n}}p_{n}^{2}+V(q)\,,\qquad m_{i}>0\,$,
is the sum of kinetic energy $T$ and a smooth homogeneous function $V(q)$ of degree $k$, then Lagrange’s identity holds
$\frac{d^{2}J}{dt^{2}}=4T-2kV=4H-2(k+2)V,\qquad J=\sum_{i=1}^{n}m_{i}q_{i}^{2}\,.$ (1.2)
Jacobi used this identity considering the stability of the motion of systems of $n$ interacting particles with masses $m_{1},\ldots,m_{n}$ arising in celestial mechanics [1], see also [2]. Various generalizations and applications of Lagrange’s identity are discussed in the works of V.V. Kozlov [3, 4].
The necessary conditions for the integrability of equations (1.1) in the class of meromorphic functions are discussed in the works [5, 6]. Below we consider Hamiltonian systems of general position, i.e., potentials that are not subject to any conditions other than the condition of homogeneity
$V(\lambda\,q_{1},\ldots,\lambda\,q_{n})=\lambda^{k}\,V(q_{1},\ldots,q_{n})\,.$
or, according to the Euler theorem for smooth homogeneous functions, the condition
$Z=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}\,V(q_{1},\ldots,q_{n})-kV(q_{1},\ldots,q_{n})=0\,$, (1.3)
which is equivalent to Lagrange’s identity (1.2) in the context of Hamiltonian mechanics.
Equations (1.1) can be interpreted geometrically in terms of the vector field $X$ on the phase space $T^{*}\mathbb{R}^{n}$, which in local coordinates $x=(q,p)$ has the form
$X=\sum_{i=1}^{n}\left(\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}-\frac{\partial V}{\partial q_{i}}\dfrac{\partial}{\partial p_{i}}\right)\,.$ (1.4)
Using the vector field $X$, Hamilton’s original equations (1.1) are rewritten in an invariant geometric form
$\dot{x}=X\,,\quad\mbox{where}\quad X=PdH\qquad\mbox{or}\qquad\iota_{X}\omega=-dH\,$, (1.5)
which contains the following basic tensor invariants: the Hamilton function $H$, the Poisson bivector $P$, and the symplectic form $\omega$, which in local coordinates have the form
$P=\sum_{i=1}^{n}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{i}}\qquad\mbox{and}\qquad\omega=\sum_{i=1}^{n}dp_{i}\wedge dq_{i}\,.$ (1.6)
The corresponding Poisson bracket on the phase space $T^{*}\mathbb{R}^{n}$ has a canonical form
$\{q_{i},p_{j}\}=\delta_{ij}\,,\qquad\{q_{i},q_{j}\}=\{p_{i},p_{j}\}=0\,,\qquad i,j=1,\ldots,n.$
In this note we prove that if potential $V(q)$ is a smooth homogeneous function of degree $k$, then for the Hamiltonian system (1.1,1.5) not only Lagrange’s identity (1.2) holds, but also exist an invariant Poisson bivector $\hat{P}$ and an invariant symplectic form $\hat{\omega}=\hat{P}^{-1}$, which cannot be obtained by differentiation and tensor operations from the basic tensor invariants $X,H,P$ and $\omega$.
1.1 Basic tensor invariants of Hamiltonian systems
Recall that smooth tensor field $T(x)$ is called a tensor invariant of system of ordinary differential equations (1.5) if
$\mathcal{L}_{X}T(x)=0\,.$ (1.7)
Here, $\mathcal{L}_{X}$ is the Lie derivative of the tensor field $T(x)$ along the vector field $X$, which determines the rate of change of the tensor field $T(x)$ during deformation of the phase space, defined by the phase flow of the system (1.5).
The vector field $X$ is a tensor invariant due to the antisymmetry of the Lie brackets $[.,.]$
$\mathcal{L}_{X}\,X=[X,X]=0\,.$
Substituting the Hamiltonian vector field $X=PdH$ into this equation, we obtain $X=PdH$
$\mathcal{L}_{X}\,X=\mathcal{L}_{X}\,\left(PdH\right)=\left(\mathcal{L}_{X}P\right)dH+Pd\left(\mathcal{L}_{X}H\right)=0\,.$ (1.8)
Both terms in this identity are equal to zero due to the definition of the Poisson bivector $P$, which is a skew-symmetric tensor field of valence $(2,0)$ (bivector)
$\mathcal{L}_{X}\,H=(dH,PdH)=0\,$,
which is invariant of the phase flow
$\mathcal{L}_{X}\,P=\mathcal{L}_{PdH}\,P=[\![P,P]\!]dH=0$ (1.9)
by virtue of the Jacobi identity
[[P,P]]=0,[\![P,P]\!]=0\,, (1.10)
which we write using the Schouten brackets [[.,.]][\![.,.]\!].
It is obvious that Jacobi’s identity (1.10) is a sufficient but not necessary condition for satisfying the invariance equation (1.9). For example, the bivector
$P^{\prime}=(a_{1}H+a_{2})P\,,\qquad a_{1,2}\in\mathbb{R}\,$,
does not satisfy Jacobi’s identity, but nevertheless is a tensor invariant for arbitrary values of the parameters $a_{1,2}$.
The invariance of the symplectic form is a consequence of condition $d\omega=0$ and Hamilton’s equations (1.5), since
$\mathcal{L}_{X}\omega=d(\iota_{X}\omega)+\iota_{X}(d\omega)=d(-dH)=0\,$,
which leads to the preservation of the $2n$ volume form
$\Omega=\wedge^{n}\omega=dq_{1}\wedge\cdots\wedge dq_{n}\wedge dp_{1}\cdots\wedge dp_{n}\,,\qquad\mathcal{L}_{X}\Omega=0\,.$
According to Cartan’s formula, the Lie derivative $\mathcal{L}_{X}$ acting on the differential $k$-form $\varphi$ is the anticommutator of the inner product with the outer derivative
$\mathcal{L}_{X}\varphi=\iota_{X}d\varphi+d\iota_{X}\varphi\,.$ (1.11)
Here, $\iota_{X}\varphi$ is the inner product of the vector $X$ and the form $\varphi$, that is, the (k1)(k-1)-form obtained from the form $\varphi$ after substituting the vector field $X$ for the first argument.
According to Poincaré [7] and Cartan [8], the differential form $\varphi$ is an absolute integral invariant if both terms in (1.11) are simultaneously equal to zero.
$\iota_{X}d\varphi=0\qquad\mbox{and}\qquad d\iota_{X}\varphi=0\,.$
These conditions are sufficient but not necessary for satisfying the invariance condition $\mathcal{L}_{X}\varphi=0$ with the Lie derivative given by (1.11) and, therefore, in works [7, 8], conditional integral invariants were introduced in addition to absolute invariants, see also discussion in [9, 10].
Absolute and conditional tensor invariants for some integrable and non-integrable systems have been studied in [11, 12, 13, 14]. In the next section, we consider a generalization of the results obtained in [14] to the case of arbitrary dimension of the phase space $T^{*}\mathbb{R}^{n}$.
2 Homogeneous potentials
Let us define a tensor field $P^{\prime}$ of valence $(2,0)$ that depends on the function $V(q)$ as follows
$P^{\prime}=$ $\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,.$ (2.1)
According to this definition, the tensor field $P^{\prime}$ (2.1) is a skew-symmetric field, i.e., a bivector.
Proposition 1
If $V(q)$ is a homogeneous function of degree $k$, then the tensor field $P^{\prime}$ (2.1) is a Poisson bivector that is invariant with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
Proof: The derivative of $P^{\prime}$ (2.1) along the vector field $X$ (1.4) is equal to
$\mathcal{L}_{X}P^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial Z}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial Z}{\partial q_{i}}\frac{\partial V}{\partial q_{j}}-\frac{\partial Z}{\partial q_{j}}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$,
and, therefore, for homogeneous potentials $V(q)$, the Lie derivative is equal to zero
$\mathcal{L}_{X}P^{\prime}=0\,$,
since $Z=0$ according to Euler’s theorem (1.3).
The verification of Jacobi’s identity (1.10) for arbitrary smooth functions $f,g,$ and $h$ of variables $x=(q,p)$
$\{\{f,g\},h\}+\{\{g,h\},f\}+\{\{h,f\},g\}=0$
is reduced to checking Jacobi’s conditions for the Poisson brackets between variables $x=(q,p)$, which have the form
$\{q_{i},q_{j}\}^{\prime}=$ $\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$ (2.2)
For example, let’s write out one of these Jacobi’s conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}Z=0\,$,
where $i,j,k=1,\ldots,n$. This is true because $Z=0$ according to Euler’s theorem (1.3). Similarly, Jacobi’s conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are also verified.
When $k\neq-2$, the invariant Poisson bivector $P^{\prime}$ (2.1) does not compatible with the Poisson bivector P{P} (1.6), i.e.
$P,\hat{P}]\!]\neq 0.$
In the case of a potential Jacobian field at k=2k=-2, the compatibility condition is satisfied $P,\hat{P}]\!]=0$ and there is another scalar invariant
$F=\frac{1}{2}\sum_{i,j=1}^{n}(q_{i}p_{j}-q_{j}p_{i})^{2}+2V\sum_{i=1}^{n}q_{i}^{2}\,,\qquad\mathcal{L}_{X}F=0\,.$
[…]
$P^{\prime}dH=mHX\,.$
According to [15] bivector $P^{\prime}$ is a product of invariant vector field $X=PdH$ and non-invariant Euler-type vector field $Y$ [16]
$P^{\prime}=X\wedge Y\,,\qquad Y=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}+\frac{k}{2}\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}\,$,
so that
$\mathcal{L}_{X}Y\wedge X=0\,.$
Proposition 2
If $V(q)$ is a homogeneous function of degree $k\neq\pm 2$, then the tensor field of valence $(0,2)$
$\hat{P}=\left(1-\frac{k}{2}\right)HP+P^{\prime}\,,\qquad\mbox{rank}\hat{P}=n\,$, (2.3)
is a non-degenerate Poisson bivector, and the tensor field of valence $(2,0)$
$\hat{\omega}=\hat{P}^{-1}$
is an invariant simplectic form with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
$\mathcal{L}_{X}\hat{P}=0\,,\qquad[\![\hat{P},\hat{P}]\!]=0\qquad\mbox{and}\qquad\mathcal{L}_{X}\hat{\omega}=0\,,\qquad d\hat{\omega}=0$
which is based on the fulfillment of the homogeneity condition $Z=0$ (1.3). The invariant symplectic 2-form is equal to
$\hat{\omega}=\frac{2}{(k-2)H}\left(\frac{k}{(k+2)H}dH\wedge d\alpha^{\prime}+\sum_{i=1}^{n}dq_{i}\wedge dp_{i}\right)\,$,
where 1-form $d\alpha^{\prime}$
$d\alpha^{\prime}=\sum_{i=1}^{n}p_{i}dq_{i}-\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$
is the standard Liouville form $\sum_{i=1}^{n}p_{i}dq_{i}$ on the cotangent bundle $T^{*}\mathbb{R}^{n}$ (or Poincaré 1-form or canonical form or tautological form) minus 1-form $\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$.
According to the Lie-Darboux theorem, there exists a canonical transformation
$(q^,p^)\phi$$\to(\hat{q},\hat{p})$
such that
$\hat{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\hat{q}_{i}}\wedge\frac{\partial}{\partial\hat{p}_{i}}\qquad\mbox{and}\qquad\hat{\omega}=\sum_{i=1}^{n}d\hat{p}_{i}\wedge d\hat{q_{i}}$
at least locally. This transformation can be viewed as analogous to the Bäcklund transformations for the integrable Hamiltonian systems with homogeneous potentials [17].
The following equation
$\hat{P}dH=\frac{k+2}{2}H\,PdH$
can be considered as a tensor analogue of the Lagrange identity (1.2).
It should be noted that within the framework of Galois theory of differential groups for Hamiltonian systems with homogeneous potentials [18], cases with homogeneity index $k=\pm 2$ are also of particular significance and require separate consideration.
3 Second family of potentials obtained
Using the shift $q_{i}\to q_{i}+a_{i}$ and the appropriate limit for the parameters $k$ and $a_{i}$, we obtain the equation for the new potential $U(q)$
$\bar{Z}=\sum_{i=1}^{n}a_{i}\frac{\partial}{\partial q_{i}}\,U(q_{1},\ldots,q_{n})-kU(q_{1},\ldots,q_{n})=0\,$, (3.1)
instead of Euler’s equation (1.3) and the condition
$\frac{d^{2}}{dt^{2}}\sum_{i=1}^{n}a_{i}q_{i}=kU.$ (3.2)
instead of the Lagrange condition (1.2). Let us compare a particular solution of Euler’s equation (1.3),
$V(q_{1},\ldots,q_{n})=q_{1}^{k}f\left(\frac{q_{2}}{q_{1}},\ldots,\frac{q_{n}}{q_{1}}\right)$
where $f$ is an arbitrary smooth function of projective coordinates, with a similar particular solution of equation (3.1), which has the form
$U(q_{1},\ldots,q_{n})=e^{\frac{k}{a_{1}}q_{1}}g\left(q_{2}-\frac{a_{2}}{a_{1}}\,q_{1},\ldots,q_{n}-\frac{a_{n}}{a_{1}}\,q_{1}\right)\,$,
where $g$ is an arbitrary smooth function. A special case of Hamiltonian systems with potential $U(q)$ are some of the Toda lattices, see [14].
Proposition 3
If the function $U(q)$ satisfies the equation (3.1) bivector
$\bar{P}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+a_{i}\frac{\partial U}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial U}{\partial q_{j}}-p_{j}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$, (3.3)
is a Poisson bivector invariant with respect to the phase flow of the vector field $X$
$\mathcal{L}_{X}\bar{P}^{\prime}=0\,,\qquad X=\sum_{i=1}^{n}\left(\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}-\frac{\partial U}{\partial q_{i}}\dfrac{\partial}{\partial p_{i}}\right)\,.$
Proof: The Lie derivative of the tensor field $\bar{P}^{\prime}$ (3.3) along the vector field $X$ is zero
$\mathcal{L}_{X}\bar{P}^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial\bar{Z}}{\partial q_{i}}\frac{\partial U}{\partial q_{j}}-\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}=0\,$,
since $\bar{Z}=0$ (3.1). Because of this, the tensor field $\bar{P}^{\prime}$ is a tensor invariant of $X$.
The verification of the Jacobi identity (1.10) is reduced to checking the Jacobi conditions for the Poisson brackets between variables $x=(q,p)$, which in this case take the form
$\{q_{i},q_{j}\}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}s}p_{i}p_{j}+a_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$
For example, let’s write out one of the conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}\bar{Z}=0\,$,
where $i,j,k=1,\ldots,n$, which is true since $\bar{Z}=0$ (3.1). The Jacobi conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are similarly verified.
[…]
$\bar{P}^{\prime}dH=kHX\,$,
but, as above, in the space of invariant bivectors, we can find a second non-degenerate Poisson bivector in addition to $P$ (1.6).
Proposition 4
If $k\neq 0$, then the tensor field
$\tilde{P}=-\frac{k}{2}HP+\hat{P}^{\prime}\,,\qquad\mbox{rank}\tilde{P}=n\,$, (3.4)
is a non-degenerate invariant Poisson bivector, and the tensor field
$\tilde{\omega}=\tilde{P}^{-1}$
is a symplectic form invariant with respect to the phase flow generated by the Hamiltonian vector field $X$.
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
$\mathcal{L}_{X}\tilde{P}=0\,,\qquad[\![\tilde{P},\tilde{P}]\!]=0\qquad\mbox{and}\qquad\mathcal{L}_{X}\tilde{\omega}=0\,,\qquad d\tilde{\omega}=0$
which is based on the fulfillment of the condition $\bar{Z}=0$ (3.1).
According to the Lie-Darboux theorem, there exists a canonical transformation of variables $(q~,p~)\phi$$\to(\tilde{q},\tilde{p})$ such that
$\tilde{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\tilde{q}_{i}}\wedge\frac{\partial}{\partial\tilde{p}_{i}}\qquad\mbox{and}\qquad\tilde{\omega}=\sum_{i=1}^{n}d\tilde{p}_{i}\wedge d\tilde{q_{i}}\,$.
This transformation can be viewed a counterpart of the Bäcklund transformations for integrable Hamiltonian systems.
The Poisson bivector $\tilde{P}$ satisfies the relation
$\tilde{P}dH=\frac{k}{2}\,H\,PdH$
which includes two invariant non-degenerate Poisson tensors and can be considered as a tensor analogue of the identity (3.2) for inhomogeneous potentials.
4 Euler-type vector fields
Let us consider homogeneous Hamilton functions $H(x)$
$H(\lambda x)=\lambda^{k}H(x)$
on the phase space $M$ with coordinates $x=(x_{1},\ldots,x_{n})$ equipped with the homogeneous Poisson bivector $P(x)$
$P(\lambda x)=\lambda^{\kappa}P(x)\,$,
where $k$ and $\kappa$ could be different We suppose that bivector
$P^{\prime}=X\wedge Y\,,\qquad X=PdH\,,\qquad Y=\sum x_{i}\frac{\partial}{\partial x_{i}}$
is an invariant rank two Poisson bivector
$\mathcal{L}_{X}P^{\prime}=0\,,\qquad[\,[P^{\prime},P^{\prime}]\,]=0\,.$
Here $Y$ is the so-called Euler vector field [16], which does not the tensor invariant of Hamiltonian flow
$\mathcal{L}_{X}Y\neq 0\,,\qquad\mathcal{L}_{X}Y\wedge X=0\,.$
The sum
$\tilde{P}=P^{\prime}-\alpha(k,\kappa)HP$
is an invariant Poisson bivector at the special value of $\alpha(k,\kappa)$, rank $\tilde{P}$=rank $P$.
As an example let us consider canonical Lie-Poisson bracket on the algebra $so^{*}=so^{*}(2)\oplus so^{*}(2)$
$\{s_{i},s_{j}\}=\varepsilon_{ijk}s_{k}\qquad\mbox{and}\qquad\{t_{i},t_{j}\}=\varepsilon_{ijk}t_{k}\,,\qquad i,j,k=1,2,3.$ (4.5)
Here $\varepsilon_{ijk}$ is the Levi-Civita symbol.
Proposition 5
Associated with homogeneous quadratic Hamiltonian
$H=\sum_{i,j=1}^{6}a_{ij}x_{i}x_{j},\qquad x=(s_{1},s_{2},s_{3},t_{1},t_{2},t_{3})$
and homogeneous bivector $P$ (4.5) vector field $X=PdH$ together with the Euler vector field
$Y=\sum x_{i}\frac{\partial}{\partial x_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=X\wedge Y\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-HP$
has a rank four.
The proof consists of a direct verification.
As a second example let us consider canonical Lie-Poisson bracket on the Lie-Poisson algebra $e^{*}(3)=so^{*}(3)\ltimes{\mathbb{R}^{3}}^{*}$
$\{L_{i},L_{j}\}=\varepsilon_{ijk}L_{k},\qquad\{L_{i},p_{j}\}=\varepsilon_{ijk}p_{k},\qquad\{p_{i},p_{j}\}=0\,.$ (4.6)
For homogeneous quadratic Hamiltonian $H$ on $e^{*}(3)$ we have an analog of the previous Proposition.
Proposition 6
Associated with Hamiltonian
$H=\sum_{i,j=1}^{2}a_{ij}L_{i}L_{j}+b_{1}x_{1}+b_{2}x_{2}+b_{3}x_{3},$
and homogeneous bivector $P$ (4.6) vector field $X=PdH$ together with the Euler-type vector field
$Y=\frac{1}{2}\sum_{i=1}^{3}L_{i}\frac{\partial}{\partial L_{i}}+\sum_{i=1}^{3}p_{i}\frac{\partial}{\partial p_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=X\wedge Y\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-\frac{1}{2}HP$
has a rank four.
The proof consists of a direct verification.
5 Conclusion
According to [19, 20], integrable Hamiltonian systems possess a family of tensor invariants of valence $(2,0)$ and $(0,2)$, which cannot be obtained from the basic tensor invariants using various tensor operations; see examples in [11, 12, 13, 14].
In this paper we prove an existence of additional tensor invariants for non-integrable Hamiltonian systems satisfying the Lagrange identity on cotangent bundle $T^{*}\mathbb{R}^{n}$. We also show similar invariant Poisson bivectors for the systems on the Lie-Poisson algebras $so^{*}(4)$ and $e^{*}(3)$.
The article was prepared within the framework of the project ”International academic cooperation” HSE University.