### v3

# Homogeneous potentials, Lagrange’s identity and Poisson geometry

Abstract
The Lagrange identity expresses the second derivative of the moment of inertia of a system of material points through kinetic energy and homogeneous potential energy, from which follows the Jacobi well-known result on the instability of a system of gravitating bodies. In this work, it is proven that if a Hamiltonian system satisfies the Lagrange identity, then it possesses additional tensor invariants that are not expressed through the basic invariants existing for all Hamiltonian systems. A new class of Hamiltonian systems with inhomogeneous potentials is considered, which also possess similar additional tensor invariants.
1 Introduction
Let us consider the Hamiltonian system of differential equations
$\frac{dq_{i}}{dt}=\frac{\partial H}{\partial p_{i}}\,,\qquad\frac{dp_{i}}{dt}=-\frac{\partial H}{\partial q_{i}}\,,\qquad i=1,\ldots,n,$ (1.1)
where $q=\{q_{1},\ldots,q_{n}\}$ are coordinates in configuration space $\mathbb{R}^{n}$, $p=\{p_{1},\ldots,p_{n}\}$ are the corresponding moments in phase space $T^{*}\mathbb{R}^{n}$, and $H=H(q,p)$ is the Hamiltonfunction.
If the Hamilton function
$H=T+V=\frac{1}{2m_{1}}p_{1}^{2}+\cdots+\frac{1}{2m_{n}}p_{n}^{2}+V(q)\,,\qquad m_{i}>0\,$,
is the sum of kinetic energy $T$ and a smooth homogeneous function $V(q)$ of degree $k$, then Lagrange’s identity holds
$\frac{d^{2}J}{dt^{2}}=4T-2kV=4H-2(k+2)V,\qquad J=\sum_{i=1}^{n}m_{i}q_{i}^{2}\,.$ (1.2)
Jacobi used this identity considering the stability of the motion of systems of $n$ interacting particles with masses $m_{1},\ldots,m_{n}$ arising in celestial mechanics [1], see also [2]. Various generalizations and applications of Lagrange’s identity are discussed in the works of V.V. Kozlov [3, 4].
The necessary conditions for the integrability of equations (1.1) in the class of meromorphic functions are discussed in the works [5, 6]. Below we consider Hamiltonian systems of general position, i.e., potentials that are not subject to any conditions other than the condition of homogeneity
$V(\lambda\,q_{1},\ldots,\lambda\,q_{n})=\lambda^{k}\,V(q_{1},\ldots,q_{n})\,.$
or, according to the Euler theorem for smooth homogeneous functions, the condition
$Z=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}\,V(q_{1},\ldots,q_{n})-kV(q_{1},\ldots,q_{n})=0\,$, (1.3)
which is equivalent to Lagrange’s identity (1.2) in the context of Hamiltonian mechanics.
Equations (1.1) can be interpreted geometrically in terms of the vector field $X$ on the phase space $T^{*}\mathbb{R}^{n}$, which in local coordinates $x=(q,p)$ has the form
$X=\sum_{i=1}^{n}\left(\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}-\frac{\partial V}{\partial q_{i}}\dfrac{\partial}{\partial p_{i}}\right)\,.$ (1.4)
Using the vector field $X$, Hamilton’s original equations (1.1) are rewritten in an invariant geometric form
$\dot{x}=X\,,\quad\mbox{where}\quad X=PdH\qquad\mbox{or}\qquad\iota_{X}\omega=-dH\,$, (1.5)
which contains the following basic tensor invariants: the Hamilton function $H$, the Poisson bivector $P$, and the symplectic form $\omega$, which in local coordinates have the form
$P=\sum_{i=1}^{n}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{i}}\qquad\mbox{and}\qquad\omega=\sum_{i=1}^{n}dp_{i}\wedge dq_{i}\,.$ (1.6)
The corresponding Poisson bracket on the phase space $T^{*}\mathbb{R}^{n}$ has a canonical form
$\{q_{i},p_{j}\}=\delta_{ij}\,,\qquad\{q_{i},q_{j}\}=\{p_{i},p_{j}\}=0\,,\qquad i,j=1,\ldots,n.$
In this note we prove that if potential $V(q)$ is a smooth homogeneous function of degree $k$, then for the Hamiltonian system (1.1,1.5) not only Lagrange’s identity (1.2) holds, but also exist an invariant Poisson bivector $\hat{P}$ and an invariant symplectic form $\hat{\omega}=\hat{P}^{-1}$, which cannot be obtained by differentiation and tensor operations from the basic tensor invariants $X,H,P$ and $\omega$.
1.1 Basic tensor invariants of Hamiltonian systems
Recall that smooth tensor field $T(x)$ is called a tensor invariant of system of ordinary differential equations (1.5) if
$\mathcal{L}_{X}T(x)=0\,.$ (1.7)
Here, $\mathcal{L}_{X}$ is the Lie derivative of the tensor field $T(x)$ along the vector field $X$, which determines the rate of change of the tensor field $T(x)$ during deformation of the phase space, defined by the phase flow of the system (1.5).
The vector field $X$ is a tensor invariant due to the antisymmetry of the Lie brackets $[.,.]$
$\mathcal{L}_{X}\,X=[X,X]=0\,.$
Substituting the Hamiltonian vector field $X=PdH$ into this equation, we obtain $X=PdH$
$\mathcal{L}_{X}\,X=\mathcal{L}_{X}\,\left(PdH\right)=\left(\mathcal{L}_{X}P\right)dH+Pd\left(\mathcal{L}_{X}H\right)=0\,.$ (1.8)
Both terms in this identity are equal to zero due to the definition of the Poisson bivector $P$, which is a skew-symmetric tensor field of valence $(2,0)$ (bivector)
$\mathcal{L}_{X}\,H=(dH,PdH)=0\,$,
which is invariant of the phase flow
$\mathcal{L}_{X}\,P=\mathcal{L}_{PdH}\,P=[\![P,P]\!]dH=0$ (1.9)
by virtue of the Jacobi identity
[[P,P]]=0,[\![P,P]\!]=0\,, (1.10)
which we write using the Schouten brackets [[.,.]][\![.,.]\!].
It is obvious that Jacobi’s identity (1.10) is a sufficient but not necessary condition for satisfying the invariance equation (1.9). For example, the bivector
$P^{\prime}=(a_{1}H+a_{2})P\,,\qquad a_{1,2}\in\mathbb{R}\,$,
does not satisfy Jacobi’s identity, but nevertheless is a tensor invariant for arbitrary values of the parameters $a_{1,2}$.
The invariance of the symplectic form is a consequence of condition $d\omega=0$ and Hamilton’s equations (1.5), since
$\mathcal{L}_{X}\omega=d(\iota_{X}\omega)+\iota_{X}(d\omega)=d(-dH)=0\,$,
which leads to the preservation of the $2n$ volume form
$\Omega=\wedge^{n}\omega=dq_{1}\wedge\cdots\wedge dq_{n}\wedge dp_{1}\cdots\wedge dp_{n}\,,\qquad\mathcal{L}_{X}\Omega=0\,.$
According to Cartan’s formula, the Lie derivative $\mathcal{L}_{X}$ acting on the differential $k$-form $\varphi$ is the anticommutator of the inner product with the outer derivative
$\mathcal{L}_{X}\varphi=\iota_{X}d\varphi+d\iota_{X}\varphi\,.$ (1.11)
Here, $\iota_{X}\varphi$ is the inner product of the vector $X$ and the form $\varphi$, that is, the (k1)(k-1)-form obtained from the form $\varphi$ after substituting the vector field $X$ for the first argument.
According to Poincaré [7] and Cartan [8], the differential form $\varphi$ is an absolute integral invariant if both terms in (1.11) are simultaneously equal to zero.
$\iota_{X}d\varphi=0\qquad\mbox{and}\qquad d\iota_{X}\varphi=0\,.$
These conditions are sufficient but not necessary for satisfying the invariance condition $\mathcal{L}_{X}\varphi=0$ with the Lie derivative given by (1.11) and, therefore, in works [7, 8], conditional integral invariants were introduced in addition to absolute invariants, see also discussion in [9, 10].
Absolute and conditional tensor invariants for some integrable and non-integrable systems have been studied in [11, 12, 13, 14]. In the next section, we consider a generalization of the results obtained in [14] to the case of arbitrary dimension of the phase space $T^{*}\mathbb{R}^{n}$.
1.2 Invariant Volterra brackets
In 1895-1899 Vito Volterra proposed a formal Hamiltonian description of second-order homogeneous vector fields with two quadratic integrals of motion. Our aim is to present similar formal Hamiltonian description for homogeneous vector fields with one homogeneous integral of motion.
In [15] Volterra consider dynamical system
$\dot{x}_{k}=\sum a_{k}^{ij}x_{i}x_{j}\,,\qquad a_{k}^{ij}\in\mathbb{R}\,.$ (1.12)
with two quadratic first integrals
$F_{A}=\frac{1}{2}(Ax,x),\qquad F_{B}=\frac{1}{2}(Bx,x).$
If quadratic form $F_{A}$ is positive-definite and $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different real roots of the characteristic polynomial
$\mbox{det}(B-\lambda A)=0,$
then a linear change of variables reduces the system of differential equations (1.12) to the form
$\dot{x}_{k}=\sum_{i<j\,i,j\neq k}c^{k}_{ij}(\lambda_{i}-\lambda_{j})x_{i}x_{j}$ (1.13)
where
$c^{j}_{ki}=c^{i}_{jk}=c^{k}_{ij}=-c^{k}_{ik}=-c^{i}_{kj}=-c^{i}_{jk},$
and other $c^{k}_{ij}$ are equal to zero. The Volterra proof is based on reduction of the pair of quadratic forms $F_{A}$ and $F_{B}$ to canonical form
$F_{A}=\frac{1}{2}\left(x^{2}_{1}+\cdots+x^{2}_{n}\right),\qquad F_{B}=\frac{1}{2}\left(\lambda_{1}x^{2}_{1}+\cdots+\lambda_{n}x^{2}_{n}\right).$
Then Volterra represents equations (1.13) by using the Jacobian of the first integrals
$\dot{x}_{k}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\frac{\partial(F_{A},F_{B})}{\partial(x_{i},x_{j})}$
and by using almost Poisson bivectors
$X=P_{A}dF_{B}=P_{B}dF_{A}\,.$ (1.14)
Here bivectors $P_{A}$ and $P_{B}$ is defined by the following almost Poisson brackets between coordinates
$\{x_{i},x_{j}\}_{A}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}x_{k}\,,\qquad\{x_{i},x_{j}\}_{B}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\lambda_{k}x_{k}$
which does not satisfy the Jacobi identity in the general case.
In modern terms Volterra proposed formal Hamiltonian description (1.14) for the second order homogeneous vector field (1.12). Using the Jacobi last multiplier theory he also proved that the phase flow of the formal Hamiltonian system (1.14) preserves the volume $n$-form
$dx_{1}\wedge\cdots\wedge dx_{n}$
if all the numbers $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different. Discussion of the Volterra formal Hamiltonian description (1.14) may be found in [17, 18].
Some examples of differential equations of the form (1.14) in three-dimensional space were considered by Volterra in connection with studying the dynamics of the Earth’s poles, see collection of papers [16]. Recently these Volterra calculation was repeated for other dynamical systems in three-dimensional space [19].
Let us consider smooth vector field
$V=\sum_{i=1}^{n}V_{i}(x_{1},\ldots,x_{n})\frac{\partial}{\partial x_{i}}$
with the homogeneous components
$V_{i}(\lambda x_{1},\ldots,\lambda x_{n})=\lambda^{\alpha}V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n\,.$
The Lie bracket between $V$ and the Euler vector field $U$ is equal to
$V,U]=(\alpha-1)V\,,\qquad U=\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial x_{i}}\,.$
according to the Euler theorem
$\sum_{k=1}^{n}x_{k}\frac{\partial V_{i}(x_{1},\ldots,x_{n})}{\partial x_{k}}=\alpha V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n.$ (1.15)
Bivector
$P^{\prime}=V\wedge U$ (1.16)
is a tensor invariant of the phase flow associated with $V$ since
$\mathcal{L}_{V}\,P^{\prime}=[V,V]\wedge U+V\wedge[V,U]=0\,.$
Here $\mathcal{L}_{V}\,P^{\prime}$ is a Lie derivative of the tensor field $P^{\prime}$ along the vector field $V$. Moreover, bivector $P^{\prime}$ (1.16) satisfies to the Jacobi condition since
$P^{\prime},P^{\prime}]\!]=2V\wedge[U,V]\wedge U=0\,$,
where [[.,.]][\![.,.]\!] is a Schouten-Nijenhuis bracket.
Let us suppose that function $F(x)$ is a homogeneous first integral of the vector field $V$, i.e.
$V(F)=\sum_{i=1}^{n}V_{i}\frac{\partial F}{\partial x_{i}}=0\qquad\mbox{and}\qquad U(F)=\sum_{i=1}^{n}x_{i}\frac{\partial F}{\partial x_{i}}=\kappa F\,.$
In this case we have
$P^{\prime}dF=-V(F)\,U+U(F)\,V=\kappa F\,V\,$,
that allows us to represent homogeneous vector field $V$ in the formal Hamiltonian form
$V=P^{\prime}dH\,,\qquad H=\kappa^{-1}\ln|F|\,.$ (1.17)
Here $P^{\prime}$ is the rank two Poisson bivector (1.16) and $H$ is a Hamilton function.
We suppose that if $V=PdH$ is a Hamiltonian vector field defined by homogeneous $P$ and $H$, the sum
$\tilde{P}=P^{\prime}-\alpha(k,\kappa)HP$
is an invariant Poisson bivector at the special value of $\alpha(k,\kappa)$, rank $\tilde{P}$=rank $P$, i.e.
$\tilde{P},\tilde{P}]\!]=2\alpha[\![P^{\prime},HP]\!]+\alpha^{2}[\![HP,HP]\!]=2\alpha[\![PdH\wedge U,HP]\!]+\alpha^{2}[\![HP,HP]\!]=0\,.$
As an example let us consider canonical Lie-Poisson bracket on the algebra $so^{*}=so^{*}(2)\oplus so^{*}(2)$
$\{s_{i},s_{j}\}=\varepsilon_{ijk}s_{k}\qquad\mbox{and}\qquad\{t_{i},t_{j}\}=\varepsilon_{ijk}t_{k}\,,\qquad i,j,k=1,2,3,$ (1.18)
and the corresponding homogeneous bivector $P_{s}$. Here $\varepsilon_{ijk}$ is the Levi-Civita symbol.
Proposition 1
Associated with homogeneous quadratic Hamiltonian
$H=\sum_{i,j=1}^{6}a_{ij}x_{i}x_{j},\qquad x=(s_{1},s_{2},s_{3},t_{1},t_{2},t_{3})$
and homogeneous bivector $P_{s}$ (1.18) vector field $V=P_{s}dH$ together with the Euler vector field
$U=\sum x_{i}\frac{\partial}{\partial x_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-HP$
has a rank four.
The proof consists of a direct verification.
As a second example let us consider canonical Lie-Poisson bracket on the Lie-Poisson algebra $e^{*}(3)=so^{*}(3)\ltimes{\mathbb{R}^{3}}^{*}$
$\{L_{i},L_{j}\}=\varepsilon_{ijk}L_{k},\qquad\{L_{i},p_{j}\}=\varepsilon_{ijk}p_{k},\qquad\{p_{i},p_{j}\}=0\,$ (1.19)
and the corresponding homogeneous bivector $P_{e}$. For homogeneous quadratic Hamiltonian $H$ on $e^{*}(3)$ we have an analog of the previous Proposition which can be extended for the case of quasi-homogeneous Hamiltonian.
Proposition 2
Associated with quasi-homogeneous Hamiltonian
$H=\sum_{i,j=1}^{2}a_{ij}L_{i}L_{j}+b_{1}p_{1}+b_{2}p_{2}+b_{3}p_{3},$
and bivector $P_{e}$ (1.19) vector field $V=P_{e}dH$ together with the Euler-type vector field
$U^{\prime}=\frac{1}{2}\sum_{i=1}^{3}L_{i}\frac{\partial}{\partial L_{i}}+\sum_{i=1}^{3}p_{i}\frac{\partial}{\partial p_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U^{\prime}\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-\frac{1}{2}HP$
has a rank four.
The proof consists of a direct verification.
2 Homogeneous potentials on $T^{*}\mathbb{R}^{n}$
Let us define a tensor field $P^{\prime}$ of valence $(2,0)$ that depends on the function $V(q)$ as follows
$P^{\prime}=$ $\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,.$ (2.1)
According to this definition, the tensor field $P^{\prime}$ (2.1) is a skew-symmetric field, i.e., a bivector.
Proposition 3
If $V(q)$ is a homogeneous function of degree $k$, then the tensor field $P^{\prime}$ (2.1) is a Poisson bivector that is invariant with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
Proof: The derivative of $P^{\prime}$ (2.1) along the vector field $X$ (1.4) is equal to
$\mathcal{L}_{X}P^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial Z}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial Z}{\partial q_{i}}\frac{\partial V}{\partial q_{j}}-\frac{\partial Z}{\partial q_{j}}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$,
and, therefore, for homogeneous potentials $V(q)$, the Lie derivative is equal to zero
$\mathcal{L}_{X}P^{\prime}=0\,$,
since $Z=0$ according to Euler’s theorem (1.3).
The verification of Jacobi’s identity (1.10) for arbitrary smooth functions $f,g,$ and $h$ of variables $x=(q,p)$
$\{\{f,g\},h\}+\{\{g,h\},f\}+\{\{h,f\},g\}=0$
is reduced to checking Jacobi’s conditions for the Poisson brackets between variables $x=(q,p)$, which have the form
$\{q_{i},q_{j}\}^{\prime}=$ $\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$ (2.2)
For example, let’s write out one of these Jacobi’s conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}Z=0\,$,
where $i,j,k=1,\ldots,n$. This is true because $Z=0$ according to Euler’s theorem (1.3). Similarly, Jacobi’s conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are also verified.
When $k\neq-2$, the invariant Poisson bivector $P^{\prime}$ (2.1) does not compatible with the Poisson bivector P{P} (1.6), i.e.
$P,\hat{P}]\!]\neq 0.$
In the case of a potential Jacobian field at k=2k=-2, the compatibility condition is satisfied $P,\hat{P}]\!]=0$ and there is another scalar invariant
$F=\frac{1}{2}\sum_{i,j=1}^{n}(q_{i}p_{j}-q_{j}p_{i})^{2}+2V\sum_{i=1}^{n}q_{i}^{2}\,,\qquad\mathcal{L}_{X}F=0\,.$
[…]
$P^{\prime}dH=mHX\,.$
According to [20] bivector $P^{\prime}$ is a product of invariant vector field $X=PdH$ and non-invariant Euler-type vector field $Y$ [21]
$P^{\prime}=X\wedge Y\,,\qquad Y=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}+\frac{k}{2}\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}\,$,
so that
$\mathcal{L}_{X}Y\wedge X=0\,.$
Proposition 4
If $V(q)$ is a homogeneous function of degree $k\neq\pm 2$, then the tensor field of valence $(0,2)$
$\hat{P}=\left(1-\frac{k}{2}\right)HP+P^{\prime}\,,\qquad\mbox{rank}\hat{P}=n\,$, (2.3)
is a non-degenerate Poisson bivector, and the tensor field of valence $(2,0)$
$\hat{\omega}=\hat{P}^{-1}$
is an invariant simplectic form with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
$\mathcal{L}_{X}\hat{P}=0\,,\qquad[\![\hat{P},\hat{P}]\!]=0\qquad\mbox{and}\qquad\mathcal{L}_{X}\hat{\omega}=0\,,\qquad d\hat{\omega}=0$
which is based on the fulfillment of the homogeneity condition $Z=0$ (1.3). The invariant symplectic 2-form is equal to
$\hat{\omega}=\frac{2}{(k-2)H}\left(\frac{k}{(k+2)H}dH\wedge d\alpha^{\prime}+\sum_{i=1}^{n}dq_{i}\wedge dp_{i}\right)\,$,
where 1-form $d\alpha^{\prime}$
$d\alpha^{\prime}=\sum_{i=1}^{n}p_{i}dq_{i}-\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$
is the standard Liouville form $\sum_{i=1}^{n}p_{i}dq_{i}$ on the cotangent bundle $T^{*}\mathbb{R}^{n}$ (or Poincaré 1-form or canonical form or tautological form) minus 1-form $\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$.
According to the Lie-Darboux theorem, there exists a canonical transformation
$(q^,p^)\phi$$\to(\hat{q},\hat{p})$
such that
$\hat{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\hat{q}_{i}}\wedge\frac{\partial}{\partial\hat{p}_{i}}\qquad\mbox{and}\qquad\hat{\omega}=\sum_{i=1}^{n}d\hat{p}_{i}\wedge d\hat{q_{i}}$
at least locally. This transformation can be viewed as analogous to the Bäcklund transformations for the integrable Hamiltonian systems with homogeneous potentials [22].
The following equation
$\hat{P}dH=\frac{k+2}{2}H\,PdH$
can be considered as a tensor analogue of the Lagrange identity (1.2).
It should be noted that within the framework of Galois theory of differential groups for Hamiltonian systems with homogeneous potentials [23], cases with homogeneity index $k=\pm 2$ are also of particular significance and require separate consideration.
3 Second family of potentials obtained
Using the shift $q_{i}\to q_{i}+a_{i}$ and the appropriate limit for the parameters $k$ and $a_{i}$, we obtain the equation for the new potential $U(q)$
$\bar{Z}=\sum_{i=1}^{n}a_{i}\frac{\partial}{\partial q_{i}}\,U(q_{1},\ldots,q_{n})-kU(q_{1},\ldots,q_{n})=0\,$, (3.1)
instead of Euler’s equation (1.3) and the condition
$\frac{d^{2}}{dt^{2}}\sum_{i=1}^{n}a_{i}q_{i}=kU.$ (3.2)
instead of the Lagrange condition (1.2). Let us compare a particular solution of Euler’s equation (1.3),
$V(q_{1},\ldots,q_{n})=q_{1}^{k}f\left(\frac{q_{2}}{q_{1}},\ldots,\frac{q_{n}}{q_{1}}\right)$
where $f$ is an arbitrary smooth function of projective coordinates, with a similar particular solution of equation (3.1), which has the form
$U(q_{1},\ldots,q_{n})=e^{\frac{k}{a_{1}}q_{1}}g\left(q_{2}-\frac{a_{2}}{a_{1}}\,q_{1},\ldots,q_{n}-\frac{a_{n}}{a_{1}}\,q_{1}\right)\,$,
where $g$ is an arbitrary smooth function. A special case of Hamiltonian systems with potential $U(q)$ are some of the Toda lattices, see [14].
Proposition 5
If the function $U(q)$ satisfies the equation (3.1) bivector
$\bar{P}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+a_{i}\frac{\partial U}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial U}{\partial q_{j}}-p_{j}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$, (3.3)
is a Poisson bivector invariant with respect to the phase flow of the vector field $X$
$\mathcal{L}_{X}\bar{P}^{\prime}=0\,,\qquad X=\sum_{i=1}^{n}\left(\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}-\frac{\partial U}{\partial q_{i}}\dfrac{\partial}{\partial p_{i}}\right)\,.$
Proof: The Lie derivative of the tensor field $\bar{P}^{\prime}$ (3.3) along the vector field $X$ is zero
$\mathcal{L}_{X}\bar{P}^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial\bar{Z}}{\partial q_{i}}\frac{\partial U}{\partial q_{j}}-\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}=0\,$,
since $\bar{Z}=0$ (3.1). Because of this, the tensor field $\bar{P}^{\prime}$ is a tensor invariant of $X$.
The verification of the Jacobi identity (1.10) is reduced to checking the Jacobi conditions for the Poisson brackets between variables $x=(q,p)$, which in this case take the form
$\{q_{i},q_{j}\}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}s}p_{i}p_{j}+a_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$
For example, let’s write out one of the conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}\bar{Z}=0\,$,
where $i,j,k=1,\ldots,n$, which is true since $\bar{Z}=0$ (3.1). The Jacobi conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are similarly verified.
[…]
$\bar{P}^{\prime}dH=kHX\,$,
but, as above, in the space of invariant bivectors, we can find a second non-degenerate Poisson bivector in addition to $P$ (1.6).
Proposition 6
If $k\neq 0$, then the tensor field
$\tilde{P}=-\frac{k}{2}HP+\hat{P}^{\prime}\,,\qquad\mbox{rank}\tilde{P}=n\,$, (3.4)
is a non-degenerate invariant Poisson bivector, and the tensor field
$\tilde{\omega}=\tilde{P}^{-1}$
is a symplectic form invariant with respect to the phase flow generated by the Hamiltonian vector field $X$.
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
$\mathcal{L}_{X}\tilde{P}=0\,,\qquad[\![\tilde{P},\tilde{P}]\!]=0\qquad\mbox{and}\qquad\mathcal{L}_{X}\tilde{\omega}=0\,,\qquad d\tilde{\omega}=0$
which is based on the fulfillment of the condition $\bar{Z}=0$ (3.1).
According to the Lie-Darboux theorem, there exists a canonical transformation of variables $(q~,p~)\phi$$\to(\tilde{q},\tilde{p})$ such that
$\tilde{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\tilde{q}_{i}}\wedge\frac{\partial}{\partial\tilde{p}_{i}}\qquad\mbox{and}\qquad\tilde{\omega}=\sum_{i=1}^{n}d\tilde{p}_{i}\wedge d\tilde{q_{i}}\,$.
This transformation can be viewed a counterpart of the Bäcklund transformations for integrable Hamiltonian systems.
The Poisson bivector $\tilde{P}$ satisfies the relation
$\tilde{P}dH=\frac{k}{2}\,H\,PdH$
which includes two invariant non-degenerate Poisson tensors and can be considered as a tensor analogue of the identity (3.2) for inhomogeneous potentials.
4 Conclusion
According to [24, 25], integrable Hamiltonian systems possess a family of tensor invariants of valence $(2,0)$ and $(0,2)$, which cannot be obtained from the basic tensor invariants using various tensor operations; see examples in [11, 12, 13, 14].
In this paper we prove an existence of additional tensor invariants for non-integrable Hamiltonian systems satisfying the Lagrange identity on cotangent bundle $T^{*}\mathbb{R}^{n}$. We also show similar invariant Poisson bivectors for the systems on the Lie-Poisson algebras $so^{*}(4)$ and $e^{*}(3)$.
The article was prepared within the framework of the project ”International academic cooperation” HSE University.